OpenStudy (anonymous):

Eliminate the “z” variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1

OpenStudy (anonymous):

2x-y+2z=-15 2x-2y-z=1 2x -y + 2z= -15 2(2x-2y-z)= 2(1) which is 2x -y + 2z= -15 4x-4y -2z = 2 You now add the like components of the equation (i.e. x with x, y with y, and constants with constants). This gives you: (2x+4x)+(-y-4y)+(2z-2z)=(-15+2) which simplifies to: 6x-5y+0=-13